Depreciation by Three Methods; Partial Years

Perdue Company purchased equipment on April 1 for $35,100. The equipment was expected to have a useful life of three years, or 7,560 operating hours, and a residual value of $1,080. The equipment was used for 1,400 hours during Year 1, 2,600 hours in Year 2, 2,300 hours in Year 3, and 1,260 hours in Year 4.

Required:

Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) the units-of-activity method, and (c) the double-declining-balance method.

Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar.

a. Straight-line method

Year Amount
Year 1 $
Year 2 $
Year 3 $
Year 4 $

b. Units-of-activity method

Year Amount
Year 1 $
Year 2 $
Year 3 $
Year 4 $

c. Double-declining-balance method

Year Amount
Year 1 $
Year 2 $
Year 3 $
Year 4 $

Please answer in the format I have provided,

Thanks
Answer #1

a. Straight-line method

Year Amount
Year 1 (35100-1080/3*9/12) = 8505
Year 2 $11340
Year 3 $11340
Year 4 $2835

b. Units-of-activity method

Year Amount
Year 1 $6300
Year 2 $11700
Year 3 $10350
Year 4 $5670

c. Double-declining-balance method

Year Amount
Year 1 $23400*9/12 = 17550
Year 2 $23400*3/12+11700*9/12 = 14625
Year 3 $1845
Year 4 $0